and 0 °C respectively. Determine the steady state heat rate through this rod via conduction. P6.4 Determine the amount of heat absorbed by a 750g piece of Cu so that it can go from 10 oc to 40 oC. A black body object whose surface area is 30 cm2 is at a temperature of 200 °K. Determine the power it emits via radiation. P6.5p6.4
Answer #1

p6.4

Specific heat of copper, C = 0.0932 cal / ( g.Co )

    Mass, m = 750 g

    Raise of temperature, ΔT = 40 - 10 = 30Co

    Heat absorbed, Q = m C ΔT

                                = 750 * 0.0932 * 30

                                = 2097 cal

                                = 2097 * 4.186 J

                                = 8778 J